PART ONE To be solved on sheet #1 EXERCISE 1 You are the controller of WTF Co. and you decide to experiment by applying the principles of ABC to the two products currently made and sold by your company. Details of the two products and some relevant information are given below for one period: Product A Product B 180 200 240 200 Output sold (in units) Output produced (in units) Costs per unit: Direct material Direct labor Machine hours (per unit) Number of components (per unit) $70 $ 110 $28$21 15 The two products are similar and are usually produced in production runs of 20 units and sold in batches of 10 units. The production overhead is currently absorbed by using a machine hour rate, and the total of the production overhead for the period has been analyzed as follows: Machine department costs (rent, business rates, depreciation and supervision) 12,430 Set-up costs Inspection/Quality control Materials handling and dispatch 5,250 4,700 7,620 You have ascertained that the cost drivers to be used are as listed below for the overhead costs shown: Cost Set up cost:s Inspection/Quality control Materials handling and dispatch Cost driver Number of production runs Number components Orders executed The number of orders executed was 42, each order being for a batch of 10 units. You are required:

a) to calculate the total costs for each product if all overhead costs are absorbed on a machine hour basics

b) to calculate the total costs for each product, using activity-based costing; to calculate and list the unit product costs from your figures in (a) and (b) above, to show the differeneces and to comment briefly on any conclusions which may be drawn which could have pricing and profit implications.
Answer #1

a)Calculation of total cost if overhead cost absorbed on a machine hour basis:

Product A

Product B

Direct Material

14,000

22,000

Direct Labor

5,600

4,200

Overheads

8,182

21,818

Total Cost

27,782

48,018

Number of Units

200

200

Cost per Unit

138.91

240.09

Machine hours of Product A = 200*3 = 600

Product B = 200*8 = 1,600

Overheads Product A = 30,000*600/2200 = $8,182

b)Using ABC:

Cost pool

Overhead

Amount

Basis

Quantity

Rate/Activity

Machine Dept Cost

12,430

Machine Hours

2,200

5.65

Set-up Cost

5,250

Number of Production Runs

20

262.5

Inspection/QC

4,700

Number of Components

4,400

1.07

Material Handling and Dispatch

7,620

Orders Executed

42

181.43

Calculation of Cost:

Product A

Product B

Direct Material

14,000

22,000

Direct Labor

5,600

4,200

Overheads:

Machine Department Cost

3,390

9,040

Set up Cost

2,625

2,625

Inspection

3,205

1,495

Material Handling and Dispatch

4,354

3,266

Total Cost

33,174

42,626

Number of Units

200

200

Cost per Unit

165.87

213.13

Comparison:

Product A

Product B

As per Traditional System

138.91

240.09

As per ABC

165.87

213.13

Difference

(26.96)

26.96

Product A will show lower profits as per traditional system i.e. overheads on the bases of machine hours

ABC is a better system of allocation of overhead cost